OpenStudy (anonymous):

Use the Law of Levers to solve the problem, x1y1 = x2y2. A rectangle has a length of 50 inches and a width of 20 inches. What is the length of another rectangle of the same area, if it has a width of 25 inches?

OpenStudy (gabylovesyou):

50*20=25*x 2 * 20= x 40=x

OpenStudy (gabylovesyou):

Do you have 40 in your answers??

OpenStudy (anonymous):

I do!

OpenStudy (gabylovesyou):

ok ;)